OpenStudy (astrophysics):

Been doing math all day, kind of brain fried, so can someone explain to me why this is? (trippple integral)

OpenStudy (astrophysics):

\[\int\limits \int\limits \int\limits (rcos \theta+rsin \theta+z)r dz dr d \theta \implies \int\limits \int\limits \int\limits (r^2 \cos \theta + r^2 \sin \theta + r z ) dz dr d \theta\] ignoring bounds nvm just figured it out xD

OpenStudy (astrophysics):

Astro you genius you.

OpenStudy (kmeis002):

gotta love that distributive property.

OpenStudy (astrophysics):

Haha it was not that :P, it was during the integration processes, one of the variables was just misplaced and typing it out I figured it out haha..

OpenStudy (kmeis002):

Ah, I see :D